The potential difference between the terminals of a battery of emf 6.0 V and internal resistance 1 Ω drops to 5.8 V when connected across an external resistor. Find emf and the internal resistance of the battery.

Grade:upto college level

1 Answers

Navjyot Kalra
9 years ago
Sol. E = 6 V, r = 1 Ω, V = 5.8 V, R = ? l = E/R + r = 6/R + 1, V = E – lr ⇒ 5.8 = 6 – 6/R + 1 * 1 ⇒ 6/R + 1 = 0.2 ⇒ R + 1 = 30 ⇒ R = 29 Ω.